The accountant for Whirly Manufacturing Inc. gathered the following information for the month of November:

Beginning WIP

16,000 units

Started production

100,000

Completed production

92,000

Ending WIP

24,000

The beginning inventory was 60% complete for materials and 20% complete for conversion costs. The ending inventory was 90% complete for materials and 40% complete for conversion costs.

Costs pertaining to the month of November are as follows:

Beginning inventory costs are:

Materials

$54,560

Direct labour

20,320

Factory overhead

15,240

Costs incurred during November include:

Materials

$468,000

Direct labour

182,880

Factory overhead

391,160

Using the FIFO method, the equivalent unit conversion cost for November is: (please show calculation)

Entry field with incorrect answer

$5.65

$5.83

$6.00

$6.20
Answer #1
To complete Beginning WIP 12800 =16000*(1-20%)
Units started and Completed 76000 =92000-16000
Ending WIP 9600 =24000*40%
The equivalent units conversion cost 98400
Conversion Costs incurred during November 574040 =182880+391160
Divide by Equivalent units 98400
The equivalent unit conversion cost 5.83
Option B $5.83 is correct